### v1

# Systematic Generalization in Language Models Scales with Information Entropy

Abstract
Systematic generalization remains challenging for current language models, which are known to be both sensitive to semantically similar permutations of the input and to struggle with known concepts presented in novel contexts. Although benchmarks exist for assessing compositional behavior, it is unclear how to measure the difficulty of a systematic generalization problem. In this work, we show how one aspect of systematic generalization can be described by the entropy of the distribution of component parts in the training data. We formalize a framework for measuring entropy in a sequence-to-sequence task and find that the performance of popular model architectures scales with the entropy. Our work connects systematic generalization to information efficiency, and our results indicate that success at high entropy can be achieved even without built-in priors, and that success at low entropy can serve as a target for assessing progress towards robust systematic generalization.
[…]
Systematic Generalization in Language Models Scales with Information Entropy
Sondre Wold, Lucas Georges Gabriel Charpentier, Étienne Simon Language Technology Group University of Oslo
1 Introduction
Human language is characterized by its combinatorial properties in syntax and semantics (Hadley, 1994). For example, when somebody knows what it means for Person A to see Person B, they can readily generalize to what it means for Person B to see Person A, without having seen examples of language use where Person B is used in the subject position (Fodor and Pylyshyn, 1988). This type of bidirectional systematic generalization has long been considered a key feature of human language understanding (Chomsky, 1957).
00111122223333000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111EntropyAccuracyCNNRNNTrans.
Figure 1: Systematic generalization increases with the entropy of the training data for popular sequence-to-sequence models, even without built-in priors.
Early critics of connectionism, such as Fodor and Pylyshyn (1988), argued that systematic generalization could not be implemented using continuous representations. Today, over thirty years after this critique was initially presented, researchers agree that achieving robust compositional behavior in neural networks remains an open problem, as shown by a recent survey by McCurdy et al. (2024). Indeed, research continues to show that current models struggle with systematic generalization (Lake and Baroni, 2018; Hupkes et al., 2020; Dziri et al., 2024; Wold et al., 2024; Mirzadeh et al., 2024). Although existing works have overcome some of these limitations by implementing strong priors (Gordon et al., 2019; Lake and Baroni, 2023), the reason for this apparent lack of generalization remains unclear: are limitations caused by inherent architectural constraints, or due to distributional properties of the data used to train models?
[…]
Figure 2: A schematic overview of the two approaches to increasing the entropy level $\mathrm{H}$. Top: Vertical scaling by distribution mixing. Bottom: Horizontal scaling is done by incrementing the support of the distribution.
In this paper, we ask the following question: Can neural networks display systematic out-of-distribution generalization without any built-in priors? To answer this, we take a data-centric approach. Specifically, we show how common architectures’ degree of systematic generalization relates to the information entropy of the distribution of component parts within training samples.
Contrary to previous work, our findings show that popular sequence-to-sequence models used in NLP are capable of systematic generalization—even without built-in architectural priors. However, this is only the case under high entropy. Through a series of sequence-to-sequence tasks based on a modified version of the SCAN (Lake and Baroni, 2018) grammar, we show that model performance scales with the entropy of the training dataset, as illustrated in Figure 1, thereby creating a link between systematic generalization and information efficiency.
In summary, our contributions are: (i) We develop and formalize a framework for studying systematic generalization based on information entropy; (ii) We experiment with two different approaches to increasing entropy, and we compare the performance of four different model architectures for these two settings; and (iii) we show that neural networks achieve systematic out-of-distribution generalization when the entropy of the training distribution is high, independently of the number of unique samples seen during training.
2 Background
Recent work by McCurdy et al. (2024, p. 9324) show that there is high agreement among researchers on the following definition of compositional behavior:
Definition 1 (CB)
“When a model receives an input I that humans conceive as composed of component parts, if the model produces correct outputs for those parts (in isolation or in other combinations), then it will also produce a correct output for I.”
This definition, however, does not consider the conditions from which a system learns to produce these outputs. Consequently, compositional behavior differs from compositional generalization. While the former can be assessed through normal benchmarks, the latter requires information about the distribution the model generalizes from.
Although generalization is a loosely defined term in machine learning research, it generally refers to how well a model performs on a held-out dataset after being trained on data that has some distributional relationship with the held-out data (Hupkes et al., 2023). When training and test distributions are similar, we refer to this as in-distribution, and when they differ, we typically refer to this as out-of-distribution. Naturally, the classification of a generalization problem as in-distribution or out-of-distribution depends on which aspects of the distribution are considered. In this work, we focus on systematic generalization as an out-of-distribution problem, where the differences in the distribution of component parts and their combinations between the training and test data are the relevant aspects.
2.1 Systematic Generalization
Following earlier works, we will refer to the systematic generalization abilities of a model as the systematicity of that model. Our view of systematicity is based on Hadley (1994), who in contrast to earlier work by Fodor and Pylyshyn (1988) distinguished between degrees of systematicity. In the following paragraphs, we define these degrees under the assumption that we are considering a system tasked to generalize from a training corpus to a held-out test corpus.
Weak systematicity.
A system displays weak systematicity if the system requires that all component parts that occur in the training data also occur at every permissible syntactic position. Even if the model succeeds at a test set that contains novel samples, none of the components in these samples occur at a novel syntactic position, essentially making the problem in-distribution with respect to the systematic properties of the data.
Quasi-systematicity.
A system displays quasi-systematicity if it (i) displays weak systematicity and (ii) can process novel sentences containing embedded sentences, [^1] In this context, an embedded sentence can be thought of as a syntactic constituent in the overall sequence. such that both the overall sentence and the embedded sentences are syntactically isomorphic to a sentence in the training data, and (iii) for every successfully processed sentence, the component parts in the embedded sentence occur in the same syntactic position in a non-embedded sentence in the training-corpora. This means that a system displays quasi-systematicity if it needs to see a component part at a specific syntactic position in a non-embedded sentence if it is to generalize to that same component when used in an embedded sentence.
Strong systematicity.
Finally, a system displays strong systematicity if it (i) displays weak systematicity and (ii) it can process novel embedded sentences containing previously learned component parts that occur in syntactic positions where they do not appear in the training data, neither in embedded sentences nor in non-embedded sentences, and (iii) a significant fraction of the possible combinations of component parts and their permissible syntactic position should be left out, based on the observation that humans learn grammars based on largely incomplete data Pinker (1989).
Systematic generalization in embedded sentences.
In this work, we will target a relaxed version of strong systematicity. We relax requirement (ii) from the original definition so that we only work with samples that have embedded sentences. As we show in the following section, this makes it easier to quantify the level of systematicity required for a given generalization problem, which is the main focus of this work. Furthermore, we also relax (ii) so that it allows component parts in the test set to occur at the same syntactic position in some embedded sentence in the train set, but not in an embedded sentence that occur at the same position in the overall sequence. In order to compare models with and without compositional priors from the literature, we base our experiments on the SCAN grammar (Lake and Baroni, 2018). This grammar has few permissible syntactic positions for each part of speech, making the original requirement from Hadley (1994) too restrictive with respect to the overall sample size needed by neural networks. In the following section, we formalize our framework and elaborate on these choices.
3 Measuring Systematicity
In this work, we will demonstrate that neural networks are able to generalize to embedded sentences where component parts occur at novel syntactic positions in the overall sequence. In our experiments, we will use a modified version of the SCAN grammar. We construct datasets where each dataset sample contains an input sequence $x\in X$ and an output sequence $y\in Y$. We discuss the specifics of this dataset generation procedure, including details about the context-free grammar (CFG) that generates $x$, in Section 3.2. In this section, we define a framework for measuring the degree of systematicity displayed by a systematic generalization problem.
Throughout this paper, $x$ will comprise two embedded sentences conjoined by a conjunctive phrase. [^2] In contrast to the original SCAN, we do not have non-embedded sentences (e.g. sequences that do not have a conjunctive phrase). Consequently, $x$ can be modeled as a triple, with a conjunction $\operatorname{c}\in C$, and the two embedded sentences $e_{1},e_{2}\in\Sigma^{*}$ as arguments, $x=(e_{1},\operatorname{c},e_{2})$. The component parts of both embedded sentences are drawn from the same vocabulary $\Sigma=V\cup A$, where $V$ is a set of verbs and $A$ is a set of adverbial phrases, e.g: jump, run $\in V$, twice $\in A$, and and $\in C$. There exists a deterministic procedure $\varphi$ for transforming $x$ into $y$:
[…]
The goal of the described task is to learn a function $f$ that approximates $\varphi$, which requires modeling the conditional distribution of output sequences given input sequences in the training data, i.e. $p^{\text{train}}(y\mid x)$. See Table 1 for more examples generated by our modified SCAN grammar. Since we are interested in systematic generalization, we need to also define the distribution of component parts in $e_{1}$ and $e_{2}$, and we are specifically interested in the case where the distribution of $e_{1}$ and $e_{2}$ is different in the training and test data, but they share a similar distribution when marginalized over position, which in this case refers to which side of the conjunctive phrase they appear in.
In SCAN (Appendix A), $e_{1}$ and $e_{2}$ always contain one verb in $V$ and zero or more adverbial phrases from $A$ that control how many times that verb should be repeated, and in what direction. To create a systematic generalization problem that adheres to the definition given in Section 2.1, it is necessary to set the support of $e_{1}$ and $e_{2}$ to be different in the training and test distributions. For simplicity, we limit our study to focus on the setting where the support of the verbs $V$ is different between the two distributions. This means that we can create a generalization problem by restricting a verb $v_{1}$ to never occur in $e_{1}$ but to occur in $e_{2}$ for $p^{train}$, as illustrated by Figure 3.
For $p^{\text{test}}$ we invert the distribution of $e_{1}$, making it a degenerate distribution with single-point support for the verb that is not in $e_{1}$ for $p^{\text{train}}$, while $e_{2}$ has the same distribution as in $p^{\text{train}}$. When generalizing from $p^{\text{train}}$ to $p^{\text{test}}$ in this scenario, a model will be exposed to all verbs in all permissible syntactic positions in some embedded sequence, but $v_{1}$ will never be seen in $e_{1}$ during training, while $e_{1}$ in $p^{\text{test}}$ always contains $v_{1}$.
[…]
Figure 3: Example of distributions of verbs for $e_{1}$ (left) and $e_{2}$ (right).
The main contribution of this work is to demonstrate that the difficulty of generalizing from $p^{\text{train}}$ to $p^{\text{test}}$ in this scenario depends on the entropy $\mathrm{H}$ of the distribution of verbs in $e_{2}$ for $p^{\text{train}}$:
$\mathrm{H}^{\text{train}}_{e_{2}}(\textrm{V})=-\sum_{v\in V}p^{\text{train}}_{e_{2}}(v)\log_{2}p^{\text{train}}_{e_{2}}(v),$
where V is the verb random variable and $V$ is the set of all verbs. For brevity, we will refer to $\mathrm{H}^{\text{train}}_{e_{2}}(\textrm{V})$ as $\mathrm{H}$. When the verbs in $e_{2}$ have a degenerate distribution, as shown in Figure 4, the entropy is zero, and when $e_{2}$ is distributed uniformly, as shown in Figure 3, the entropy is at its maximum. Decreasing $\mathrm{H}$ within this interval increases the difficulty of the systematic generalization problem. Note that even at maximum entropy, the generalization problem is still out-of-distribution, as the restricted verb is never seen in $e_{1}$. In Section 4, we show empirically that the performance of neural network architectures scales with $\mathrm{H}$ on this type of generalization problem.
The degenerate case.
[…]
Figure 4: Probabilities in the degenerate case for $e_{2}$
We note that $\mathrm{H}=0$ constitutes a special case in the described framework. If $e_{2}$ is described by a degenerate distribution, then instances of $e_{2}$ will always contain the same $v_{1}\in V$. Now, if $|C|=1$, there is no way of decoupling the semantics of the conjunctive phrase and $v_{1}$, as these will always appear together in $X$. In this case, we argue, no system will be able to generalize from $p^{\text{train}}$ to $p^{\text{test}}$. Consequently, it is necessary to set $|C|>1$ for evaluating performance at $\mathrm{H}=0$.
Input sequence $x$ Output sequence $y$
squat opposite right and squat RTURN RTURN SQUAT SQUAT
squat twice and crawl opposite left SQUAT SQUAT LTURN LTURN CRAWL
sprint right twice after sprint left LTURN SPRINT RTURN SPRINT RTURN SPRINT
lunge opposite left and look thrice LTURN LTURN LUNGE LOOK LOOK LOOK
Table 1: Examples from our generated datasets, with green indicating $e_{1}$ and orange indicating $e_{2}$.
3.1 Increasing Entropy
In this work, we experiment with two different approaches to increasing $H$, as illustrated in Figure 2. In the following sections we formalize these approaches, and in Section 4 we demonstrate how the performance of neural architectures scales for each type of increment.
3.1.1 Distribution Mixing
Let $\mathcal{U}$ be a uniform distribution with support over $V\setminus\{v_{1}\}$, and let $\mathcal{D}$ be a degenerate distribution over $V$ with single-point support for $v_{1}$. Our first approach consists in defining a family of distributions parametrized by $\lambda$ mixing $\mathcal{U}$ and $\mathcal{D}$:
$p_{e_{2}}^{\text{train}}(\mathrm{V)}=\lambda\mathcal{U}+(1-\lambda)\mathcal{D},$
Then, the entropy $\mathrm{H}$ is at the minimum when $\lambda=0$ and at the maximum when $\lambda=1-1/|V|$, which is equal to a uniform distribution over $V$. Consequently, increasing $\mathrm{H}$ can be achieved by tending $\lambda$ towards $1-1/|V|$. This approach to increasing $\mathrm{H}$ is also illustrated in the top part of Figure 2.
3.1.2 Incremental Support
In our second approach, we define another family of distributions, all uniforms over increasingly larger supports $S_{i}\subseteq V$:
$\{v_{1}\}=S_{1}\subset S_{2}\subset\cdots\subset S_{|V|}=V$
The uniform distribution over the smallest support $S_{1}$ is equivalent to the degenerate distribution $\mathcal{D}$. As $i$ increases, $\mathrm{H}$ increases logarithmically until reaching maximum entropy at $i=|V|$ where we have the uniform distribution over $V$. This approach to increasing $\mathrm{H}$ is also illustrated in the bottom part of Figure 2.
3.2 Data Generation
As posited in the previous sections, the quantification of systematic generalization requires information about the distribution of parts and combinations observed during training. This level of granularity is generally intractable when dealing with real-world corpora. Without this information, however, it is difficult to control for statistical patterns that permit a non-compositional solution. Studies on systematicity typically overcome this problem by studying synthetic languages where the distribution can be known a priori. This is also the approach taken in this work.
To generate datasets samples $(x,y)\in X\times Y$ we define a context-free grammar (CFG) based on the SCAN grammar (Lake and Baroni, 2018). An overview of the complete vocabulary can be found in Appendix A, while Table 1 shows examples from our generated data. Importantly, we set $|V|$ to 8, as opposed to the original four verbs, thereby increasing the number of possible sequences generated by the grammar. We also remove the turn operators, as they are more syntactically constrained than other verbs, as was already done in Gordon et al. (2019). As previously stated, all instances of $x$ contain embedded sentences $e_{1}$ and $e_{2}$.
In our experiments, we use the standard conjunctive phrases in SCAN $|C|=2$. Furthermore, we let the two conjunctive phrases take the embedded sentences in the opposite order: if $v_{1}$ always occur in $e_{2}$ for $c_{1}$, then $v_{1}$ will always occur in $e_{1}$ for $c_{2}$.
4 Experiments
We validate our framework by evaluating the three most common sequence-to-sequence architectures: the Transformer, the RNN, and the CNN. We also run a model with built-in compositional priors to showcase how a model that encodes verb equivariance can solve the task even when $\mathrm{H}$ is low. We have two main experimental settings, corresponding to the two methods for increasing $\mathrm{H}$ outlined in Section 3.1.1 and Section 3.1.2. We also conduct two supporting experiments. For all experimental settings, $\mathrm{H}=0$ constitutes the most difficult generalization problem, and $\mathrm{H}=\log_{2}|V|=3$ constitutes the easiest problem. We average and report results from five differently seeded runs. The following sections provide details about our experimental setup. [^3] All code related to our experiments will be released upon publication at anonymous-url.
4.1 Models
RNN.
For our RNN we use a similar configuration as previous work on SCAN (Lake and Baroni, 2018; Gordon et al., 2019): a bidirectional encoder-decoder with attention. Details on the implementation and choice of hyperparameters can be found in Section B.1.
CNN.
For our CNN we use the encoder-decoder architecture from Gehring et al. (2017), which has been proven effective on SCAN in previous work (Dessì and Baroni, 2019). As in the RNN, this CNN uses an attention mechanism between the encoder and decoder. Hyperparameters can be found in Section B.2.
Transformer.
We use the original encoder-decoder Transformer from Vaswani (2017). We use this for two reasons. Firstly, this makes comparison with previous work on systematicity easier, as this has been used in previous work such as Hupkes et al. (2020) and Lake and Baroni (2023). Secondly, recent work on machine translation has shown that encoder-decoders outperform the more widely used decoder-only formulation of the Transformer on similar parameter sizes (Pitorro et al., 2024). As SCAN is framed and modeled as a sequence-to-sequence task, the encoder-decoder formulation is the most reasonable choice. We use absolute position encodings for the two main experiments, as well as Gated Linear Units as the activation function (Shazeer, 2020). Details on the implementation and choice of hyperparameters can be found in Section B.3.
Permutation-equivariant model.
Lastly, we experiment with a model that has built-in architectural priors. We use the encoder-decoder from Gordon et al. (2019), which enforces the verb equivariance found in SCAN using a cyclic group. This group is used in the encoding of the input sequences in a way that pools the representation of any verb with the representation of the other verbs, even when they are not present in the input. This is a strong structural prior that makes the model able to solve the original SCAN splits efficiently. However, this approach is dataset-dependent, as the elements of the cyclic group must be defined manually. We use this model primarily as a sanity check for our experiments, but also as a comparison for the other architectures, comparing their efficiency to an idealized case.
4.2 Experiment 1: Vertical scaling
For Experiment 1, $p^{\text{test}}$ has a degenerate distribution for $e_{1}$, where $v_{1}\in V$ always occur, while $e_{2}$ has a uniform distribution over $V$. For $p^{\text{train}}$, $e_{1}$ has a uniform distribution over $V\setminus\{v_{1}\}$, and $e_{2}$ is a mixture of a degenerate distribution at $v_{1}$ and a uniform distribution over $V\setminus\{v_{1}\}$, where we increase the total samples drawn from the uniform while decreasing the number drawn from the degenerate distribution. By doing this we increase $\mathrm{H}$ as described in section Section 3.1.1. We refer to this increase in $\mathrm{H}$ as vertical scaling, as the distribution over the verbs increases towards the uniform. The models are trained on 6 000 unique samples and evaluated on 7 056 samples for all values of $\mathrm{H}$.
00111122223333000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111EntropyAccuracyCNNRNNTrans.PermEq.
(a) Experiment 1: Vertical scaling of entropy
00111122223333000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111EntropyAccuracyCNNRNNTrans.PermEq.
(b) Experiment 2: Horizontal scaling of entropy
Figure 5: Results from the experiment 1 & 2. The accuracy and standard deviations are from five seeds.
4.3 Experiment 2: Horizontal Scaling
For Experiment 2, $p^{\text{test}}$ is the same as for Experiment 1. For $p^{\text{train}}$, $e_{1}$ is also the same as for Experiment 1, but $e_{2}$ is a uniform distribution over an increasing number of verbs, starting with $\{v_{1}\}$ and ending at $V$, as described in Section 3.1.2. We refer to this increase in $\mathrm{H}$ as horizontal scaling, as the increase of $\mathrm{H}$ is achieved by adding another verb to the support of $e_{2}$. The models are trained on all grammatical combinations in $\Sigma^{*}$ permitted by the support at each level of $\mathrm{H}$. When $\mathrm{H}=2$, for example, the support of the distribution of $e_{2}$ contains four verbs, and the training set contains all possible pairs $(x,y)\in X\times Y$ under these constraints.
To ensure a fair comparison across entropy levels, we train all models for the same number of optimization steps, so for $\mathrm{H}=1$, which has fewer training samples than $\mathrm{H}=3$, we do more iterations over the full data.
4.4 Effect of Number of Unique Samples
As $H$ describes the relative distribution of verbs in $e_{2}$, it is possible to construct datasets of varying sizes that have the same $\mathrm{H}$. This can be achieved by increasing the total volume of unique samples while maintaining the same relative distribution. To verify that performance scales with $\mathrm{H}$ independently of the number of unique samples seen during training, we create datasets that have the same entropy as in the previous experiments, but where we vary the number of total unique samples. We experiment with three different sample sizes: 3 000, 4 000, and 6 000, using the vertical scaling setting. All models run for the same number of optimization steps for all sample sizes.
4.5 Effect of Positional Embedding Type
In our experiments, there is an intuitive connection between the generalization from $p^{\text{train}}$ to $p^{\text{test}}$ and the ability to encode verb equivariance. As the verbs always occur on the first token position of $e_{1}$ and $e_{2}$, the positional information is central to both of our experiments.
To assess the effect of positional information, we show how the choice of positional encoding scheme affects performance in the vertical scaling setting. In addition to the original absolute position embeddings from Vaswani (2017), we experiment with embeddings that encode relative positional information through disentangled attention (He et al., 2021) and rotary position embeddings (RoPE; Su et al., 2024). In this experiment, we focus on the Transformer. Details on hyperparameter selection can be found in Section B.4
4.6 Results and Discussion
The results of Experiment 1 and Experiment 2 can be found in Figure 5. We observe a clear positive relationship between performance and $\mathrm{H}$ for both experiments, demonstrating that systematic generalization scales with entropy. Unsurprisingly, the permutation-equivariant model, which has a strict compositional prior, solves both tasks at all entropy levels. This shows that the task is solvable even at $\mathrm{H}=0$ when enforcing such a prior, but that vanilla architectures generally require much higher entropies for this type of generalization.
In Experiment 1, the Transformer outperforms both the RNN and CNN across all levels of $\mathrm{H}$. In Experiment 2, the performances of all models are statistically similar, falling within one standard deviation of each other, but with the CNN performing the best on average in the reported runs.
We also observe that for the Transformer, a reduction in the support of the distribution is more detrimental to the performance than a step towards the degenerate distribution: At $\mathrm{H}=2$, the Transformer has close to 100% accuracy for Experiment 1, but for the same level of $\mathrm{H}$ in Experiment 2, it achieves around 80% accuracy. This is not the case for the RNN and CNN, where we observe the opposite: The increase in support that increases $\mathrm{H}$ from 1 to 2 in Experiment 2 gives a higher performance increase than what the same increase in $\mathrm{H}$ does for Experiment 1.
We argue that these results indicate that the Transformer is more information efficient than the RNN and CNN under full support, but that the models are similar under lacking support. In practice, this means that the Transformer is able to generalize to novel uses of component parts, as long as the overall syntactic pattern—constituted by the conjunctive phrase in our experiments—is frequent in the training data. This interpretation aligns with Loula et al. (2018), showcasing how our framework can provide a theoretical explanation for previous results.
We note that even though all models can generalize under high entropy for both experiments, their performance is subpar for low entropy, e.g. $\mathrm{H=1}$. This indicates a rather high lower bound on the entropy required for systematic generalization, both for vertical and horizontal scaling. This shows that systematic generalization remains an open problem when there is no strict architectural prior, and it remains unclear whether or not it is possible to build a model that performs well on low entropy without such priors.
4.6.1 Supplemental results
00111122223333000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111EntropyAccuracy300040006000
(a) Effect of unique sample size on the Transformer
00111122223333000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111EntropyAccuracyRotaryRelativeAbsol.
(b) Effect of positional embedding type for the Transformer
Figure 6: The result of the supplemental experiments. The accuracy and standard deviations are from five seeds.
In Figure 6(a), we show the results of varying the number of unique samples at each level of $\mathrm{H}$. We observe that the positive relationship between performance and $\mathrm{H}$ is independent of the number of unique samples. The Transformer performs similarly when trained on 3 000 to 6 000 unique samples. During hyperparameter selection (Appendix B), we also observe that performance does not scale with model size. This indicates that achieving higher accuracy on lower levels of $\mathrm{H}$ is not achievable by scaling the number of unique data samples or model size for the same architecture.
In Figure 6(b), we report the performance of different positional encoding types. Although all encoding types achieve near-ceiling performance at high entropy, absolute encoding performs better for all levels of $\mathrm{H}$. From our task formulation in Section 3, it is clear that positional information is crucial for solving the tasks. We attribute the efficiency of absolute encodings to the fact that the positional information of the verbs is independent of the remainder of the sequence. This invariance has to be learned by the other embedding types, making the absolute variant more information efficient.
5 Previous Work
In this work, we base our experiments on a version of the SCAN grammar from Lake and Baroni (2018). The original grammar has been studied in other works, such as Loula et al. (2018); Dessì and Baroni (2019); Gordon et al. (2019), and there has also been work on the related PCFG SET dataset from Hupkes et al. (2020).
Most similar to our work is Zhou et al. (2023), who also studies how aspects of the SCAN grammar affects compositional generalization. We see compositionality as a set of primitives composed by a composition function, and our work focuses on how models capture this composition function depending on properties of the primitives. Their work, however, do not make this distinction, as their complexification process modify both the composition function and the distribution of primitives. The main distinction is hence that our work relates compositionality to dataset-level regularities, while they relate it to sample-level features.
In another domain, Wiedemer et al. (2023) approaches systematicity in an image generation problem by factorizing the latent variables of the input image. The authors show that the support of these variables must be the same in the training and test distributions for any model to generalize systematically. We take a more fine-grained approach, and our results demonstrate that support alone is insufficient as a predictor of systematic generalization; the performance also depends on the entropy, and higher levels of entropy in the training distribution can compensate for lower support.
There is little existing research on how to measure systematicity from a data-centric perspective. However, there is ongoing work on formalizing compositionality in terms of data complexity (Elmoznino et al., 2024) and kernel-theory (Lippl and Stachenfeld, 2024). Most similar to our approach is Keysers et al. (2020), which also quantifies the level of systematic generalization using distributional information. Specifically, they quantify the compositionality of an experiment as the divergence between the distribution of component parts and combinations in the training and test data. Their work differs from ours since they focus on the case where the distribution of component parts is as similar as possible between the training and test data, while our work focuses on the case where the distribution of component parts is inverted from the training to the test set for some parts of the input.
6 Conclusion
In this work, we show that the performance of sequence-to-sequence models scales with entropy. We also show that this scaling is independent of the number of unique samples. Contrary to previous work, our findings demonstrate that these models are capable of systematic generalization. However, this requires high entropy. This poses a new question: Is it possible to achieve systematicity also at low entropy without encoding explicit architectural priors? Through our formalizations, we hope to facilitate future work that tries to answer this question, providing a clear method for assessing progress in systematic generalization research.
Limitations
Scope of our findings.
In this work we have considered an aspect of systematicity that is concerned with embedded sentences as defined by Hadley (1994). There is no guarantee that our findings generalize to scenarios that are different from ours, but that still falls in under some definition of systematic generalization.
Use of synthetic data.
As is commonplace in compositionality research, our work relies on synthetic data that is generated by a CFG. This is necessary to quantify the entropy and to get a controlled experimental setup where the distribution of component parts in the embedded sentences can be measured precisely. This is generally not tractable for real-world corpora. Consequently, our method is constrained to cases where fine-grained distributional information is obtainable.
Appendix A Modified SCAN grammar
$V$ $A$
look left
jump right
run opposite left
walk opposite right
sprint around left
crawl around right
squat twice
lunge thrice
Table 2: The verbs and adjective phrases in our modified version of SCAN.
Our modified SCAN grammar consists of the same parts of speech as the original, but we extend the set of verbs $V$ with four additional items. We keep the original eight adjective phrases. The verbs and adjective phrases can be seen in Table 2. Every verb can be combined with every adjective phrase in each embedded sequence (under no distributional constraints), combined with one of the two conjunctive phrases from SCAN (and, after).
Appendix B Model details
All models were implemented in Pytorch (Paszke et al., 2019). We conducted hyperparameter searches for each model per experiment. Below is a list of the search space used for each. All configurations used a weight decay of $1\times 10^{-1}$ and a 10% dropout rate.
B.1 RNN
We use the basic sequence-to-sequence implementation from Gordon et al. (2019), which is based on the Pytorch MT tutorial. We conduct a grid search over the following learning rates: $\{1\times 10^{-3},1\times 10^{-4},3\times 10^{-4}\}$; cell type: Elman RNN (Elman, 1991), LSTM (Hochreiter and Schmidhuber, 1997), GRU (Chung et al., 2014); number of hidden units: $\{64,128,256\}$; number of layers: $\{1,2,3\}$; and teacher-forcing rate: $\{0.0,0.5\}$. All models ran with an attention mechanism between the encoder and decoder, bidirectional layers, and with no learning rate decay. Following previous work, we used a batch size of 1.
The RNN used for the experiment outlined in Section 4.2 has a learning rate of $1\times 10^{-4}$, normal RNN cell types, a hidden size of 128 and 2 layers (483 438 parameters), while for the experiment outlined in Section 4.3 it has 1 layer and 64 hidden units (112 238 parameters). Teacher-forcing did not result in increased performance for either experiment.
B.2 CNN
We based our implementation of the CNN from Gehring et al. (2017) on the one found in fairseq (Ott, 2019). We conduct a grid search over the following learning rates: $\{1\times 10^{-2},1\times 10^{-3},1\times 10^{-4},3\times 10^{-4}\}$; kernel size: $\{3,5\}$; number of hidden units: $\{64,128,256\}$; and number of layers: $\{1,2,3\}$; We train with a batch size of 32 and a cosine decay on the learning rate.
The CNN used for the experiment outlined in Section 4.2 has a learning rate of $3\times 10^{-4}$, a kernel size of 5, a hidden size of 64 and 3 layers (298 078 parameters), and the same configuration for Section 4.3.
B.3 Transformer
We implement the Transformer following the original from Vaswani (2017). We conduct a grid search over the following learning rates: $\{1\times 10^{-3},1\times 10^{-4},3\times 10^{-4},1\times 10^{-5}\}$; number of hidden units: $\{64,128,256\}$; and number of layers: $\{1,2,3\}$; As for the CNN, we train with a batch size of 32 and a cosine decay on the learning rate.
The Transformer used for the experiment outlined in Section 4.2 has a learning rate of $3\times 10^{-4}$, a hidden size of 128 and 3 layers (1 796 894 parameters), while for the experiment outlined in Section 4.3 it has 2 layers and 256 hidden units (4 763 166 parameters).
B.4 Positional Embedding
In addition to the absolute encodings used in Vaswani (2017), we implement two other variations of positional embeddings: RoPE Su et al. (2024) and Disentangled He et al. (2021). For all variations, we use the same hyperparameters found in the Transformer search (Section B.3). Additionally, we conduct a hyperparameter search for the encoding-specific parameters. For RoPE we conduct a search for the base value of the $\theta$. We tried powers of 10, from one to one million. For Disentangled Attention we sweep over the input length for the relative attention. We have a separate length for the input and output encodings. We sweep over the following values: $\{4,8,16,32,64\}$. As for the Transformer, we train with a batch size of 32 and a cosine decay on the learning rate.
For the models used for the experiment outlined in Section 4.5, we use a RoPE $\theta$ value of 1 000, and for the disentangled encodings we use a relative input length of 64 and a relative output length of 4.
B.5 Permutation equivariant sequence model
For the permutation equivariant model, we use the original implementation from Gordon et al. (2019). We extend the cyclic group from four to eight verbs. For the hyperparameter selction, we conducted a grid search over the following learning rates: $\{1\times 10^{-3},1\times 10^{-4},3\times 10^{-4},1\times 10^{-5}\}$; the number of hidden units: $\{64,128\}$; and RNN cell type: Elman RNN, GRU. All settings ran with 1 layer.
The model used for the experiment outlined in Section 4.2 has a learning rate of $1\times 10^{-4}$, a hidden size of 64 and the GRU cell types ($\approx$ 130 000 parameters), and the same configuration for Section 4.3.
We note that the model presented in Lake and Baroni (2023) also encodes verb equivariance and achieves high performance on SCAN. However, this approach relies on automatically augmenting the training data by creating examples of the input that have different semantics than the original grammar, making our quantification of $H$ impractical. Consequently, we only evaluate the permutation-equivariant model from Gordon et al. (2019).
Appendix C Computational budget
All hyperparameter sweeps and experiments ran on AMD EPYC 7763 CPUs:
  • RNN sweep: 2880 hours.
  • CNN sweep: 139 hours.
  • Transformer sweep: 120 hours.
  • Permutation equivariant sweep: 24 hours
  • RNN training run: 600 hours.
  • CNN training run: 21 hours.
  • Transformer sweep: 120 hours.
  • Evaluation: 12 hours.
  • Total: 3916 core CPU hours.
Appendix D Licenses and use
The present work uses the following scientific artifacts:
  1. 1.
    We base our synthetic data on the SCAN grammar (Lake and Baroni, 2018). SCAN is released under a BSD License. We do not redistribute or use any of the source code, nor do we redistribute or use any of the original data files. We implement the grammar based on the description provided in the paper. Our code implementation of the grammar includes the original license statement from SCAN.
  2. 2.
    Our use of the permutation equivariant model from Gordon et al. (2019) uses the original implementation. The code is released under an MIT license. All source code files redistributed with our work contain the original license statement, as per the license requirement.
  3. 3.
    The Transformer (Vaswani, 2017) and the CNN (Gehring et al., 2017) is implemented by the authors in PyTorch (Paszke et al., 2019) and do not use existing code implementations. However, the CNN was largely based on the implementation found in Fairseq (Ott, 2019), which is released under an MIT license. We do not redistribute any of the original code files.
Appendix E The use of AI assistants
The authors used AI assistants for spellchecking and sentence-level editing in the writing of the current manuscript.

### v2

@@
 Human language is characterized by its combinatorial properties in syntax and semantics (Hadley, 1994).
-For example, when somebody knows what it means for Person A to see Person B, they can readily generalize to what it means for Person B to see Person A, without having seen examples of language use where Person B is used in the subject position (Fodor and Pylyshyn, 1988).
-This type of bidirectional systematic generalization has long been considered a key feature of human language understanding (Chomsky, 1957).
+For example, when someone knows what it means for Person A to see Person B, they can readily generalize to what it means for Person B to see Person A, without having been exposed to examples where Person B is used in the subject position (Fodor and Pylyshyn, 1988).
+This type of bidirectional systematic generalization has long been considered a key feature of human language understanding (Chomsky, 1957), and is closely related to the compositionality of natural language semantics, a contentious but widely studied topic in both formal semantics (van Bethem et al., 1991) and NLP (McCurdy et al., 2024).
 00111122223333000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111EntropyAccuracyCNNRNNTrans.
 Figure 1: Systematic generalization increases with the entropy of the training data for popular sequence-to-sequence models, even without built-in priors.
-Early critics of connectionism, such as Fodor and Pylyshyn (1988), argued that systematic generalization could not be implemented using continuous representations.
-Today, over thirty years after this critique was initially presented, researchers agree that achieving robust compositional behavior in neural networks remains an open problem, as shown by a recent survey by McCurdy et al.
+Early critics of connectionism, such as Fodor and Pylyshyn (1988), argued that systematic generalization was incompatible with continuous representations.
+More than thirty years since this critique was initially presented, researchers still agree that achieving robust compositional behavior in neural networks remains an open problem, as shown by a recent survey by McCurdy et al.
 (2024).
-Indeed, research continues to show that current models struggle with systematic generalization (Lake and Baroni, 2018; Hupkes et al., 2020; Dziri et al., 2024; Wold et al., 2024; Mirzadeh et al., 2024).
-Although existing works have overcome some of these limitations by implementing strong priors (Gordon et al., 2019; Lake and Baroni, 2023), the reason for this apparent lack of generalization remains unclear: are limitations caused by inherent architectural constraints, or due to distributional properties of the data used to train models?
+Research continues to demonstrate that current models struggle with systematic generalization (Lake and Baroni, 2018; Hupkes et al., 2020; Dziri et al., 2024; Wold et al., 2024; Mirzadeh et al., 2024).
+Some of the limitations of current models can be alleviated by implementing compositional priors, either in the architecture (Gordon et al., 2019) or through the training procedure (Lake and Baroni, 2023), but it remains unclear whether architectural constraints cause the limitations or whether they arise due to distributional properties of the data used for training.
 […]
 […]
-Figure 2: A schematic overview of the two approaches to increasing the entropy level $\mathrm{H}$.
+Figure 2: A schematic overview of the two approaches to increasing the entropy level.
 Top: Vertical scaling by distribution mixing.
-Bottom: Horizontal scaling is done by incrementing the support of the distribution.
+Bottom: Horizontal scaling by incrementing the support of the distribution.
 In this paper, we ask the following question: Can neural networks display systematic out-of-distribution generalization without any built-in priors?
 To answer this, we take a data-centric approach.
-Specifically, we show how common architectures’ degree of systematic generalization relates to the information entropy of the distribution of component parts within training samples.
-Contrary to previous work, our findings show that popular sequence-to-sequence models used in NLP are capable of systematic generalization—even without built-in architectural priors.
-However, this is only the case under high entropy.
-Through a series of sequence-to-sequence tasks based on a modified version of the SCAN (Lake and Baroni, 2018) grammar, we show that model performance scales with the entropy of the training dataset, as illustrated in Figure 1, thereby creating a link between systematic generalization and information efficiency.
+Specifically, we show how the degree of systematic generalization of common architectures relates to the information entropy of the distribution of component parts found in the training data.
+Contrary to previous work, we find that popular sequence-to-sequence models used in NLP are capable of systematic generalization, even without built-in architectural priors.
+However, this is only the case when the training dataset contains high entropy.
+On a series of sequence-to-sequence datasets based on a modified version of the SCAN (Lake and Baroni, 2018) grammar, we show that model performance scales with the entropy of the training dataset, as illustrated in Figure 1, thereby creating a link between systematic generalization and information efficiency.
 In summary, our contributions are: (i) We develop and formalize a framework for studying systematic generalization based on information entropy; (ii) We experiment with two different approaches to increasing entropy, and we compare the performance of four different model architectures for these two settings; and (iii) we show that neural networks achieve systematic out-of-distribution generalization when the entropy of the training distribution is high, independently of the number of unique samples seen during training.
 2 Background
-Recent work by McCurdy et al. (2024, p. 9324) show that there is high agreement among researchers on the following definition of compositional behavior:
+Systematic generalization is typically described as a robust mode of composition, where a system is able to combine parts of its training data in a way that enables generalization to novel combinations.
+As such, systematic generalization relates to compositionality, a topic which has both a long tradition and multiple formal definitions, most prominently in formal semantics (van Bethem et al., 1991; Pagin and Westersthl, 2010).
+These definitions, however, often prove too restrictive when working with natural languages.
+As a result, researchers frequently introduce problem-specific relaxations, leading to ambiguity over time regarding the exact meaning of terms such as systematic generalization and compositionality.
+Despite there being some ambiguity regarding the formal definitions, recent work shows that there is high agreement among researchers on the following, more informal definition of what constitutes compositional behavior:
 Definition 1 (CB)
-“When a model receives an input I that humans conceive as composed of component parts, if the model produces correct outputs for those parts (in isolation or in other combinations), then it will also produce a correct output for I.”
+“When a model receives an input $I$ that humans conceive as composed of component parts, if the model produces correct outputs for those parts (in isolation or in other combinations), then it will also produce a correct output for $I$ (McCurdy et al., 2024, p.
+9324).”
 This definition, however, does not consider the conditions from which a system learns to produce these outputs.
@@
 Our view of systematicity is based on Hadley (1994), who in contrast to earlier work by Fodor and Pylyshyn (1988) distinguished between degrees of systematicity.
-In the following paragraphs, we define these degrees under the assumption that we are considering a system tasked to generalize from a training corpus to a held-out test corpus.
-Weak systematicity.
+In the following paragraphs, we present and define these degrees under the assumption that we are considering a system tasked to generalize from a training corpus to a held-out test corpus: [^1] Hadley (1994) also discusses an intermediate degree between weak and strong systematicity, so-called quasi-systematicity, but we leave this out for brevity, as it does not relate directly to our use case.
+Definition 2 (Weak systematicity)
 A system displays weak systematicity if the system requires that all component parts that occur in the training data also occur at every permissible syntactic position.
 Even if the model succeeds at a test set that contains novel samples, none of the components in these samples occur at a novel syntactic position, essentially making the problem in-distribution with respect to the systematic properties of the data.
-Quasi-systematicity.
-A system displays quasi-systematicity if it (i) displays weak systematicity and (ii) can process novel sentences containing embedded sentences, [^1] In this context, an embedded sentence can be thought of as a syntactic constituent in the overall sequence.
-such that both the overall sentence and the embedded sentences are syntactically isomorphic to a sentence in the training data, and (iii) for every successfully processed sentence, the component parts in the embedded sentence occur in the same syntactic position in a non-embedded sentence in the training-corpora.
-This means that a system displays quasi-systematicity if it needs to see a component part at a specific syntactic position in a non-embedded sentence if it is to generalize to that same component when used in an embedded sentence.
-Strong systematicity.
-Finally, a system displays strong systematicity if it (i) displays weak systematicity and (ii) it can process novel embedded sentences containing previously learned component parts that occur in syntactic positions where they do not appear in the training data, neither in embedded sentences nor in non-embedded sentences, and (iii) a significant fraction of the possible combinations of component parts and their permissible syntactic position should be left out, based on the observation that humans learn grammars based on largely incomplete data Pinker (1989).
+Definition 3 (Strong systematicity)
+Now, a system displays strong systematicity if it (i) displays weak systematicity, meaning that it can generalize in-distribution, and (ii) it can process novel embedded sentences containing previously learned component parts that occur in syntactic positions where they do not appear in the training data, neither in embedded sentences nor in non-embedded sentences, and (iii) a significant fraction of the possible combinations of component parts and their permissible syntactic position should be left out, based on the observation that humans learn grammars based on largely incomplete data Pinker (1989).
+In this context, an embedded sentence can be thought of as a syntactic constituent in the overall sentence.
+For example, if a system can process the following novel sentence Mathilde knows Sigurd and Sigurd knows Olav, which has the sentences Mathilde knows Sigurd and Sigurd knows Olav as embedded sentences, without having seen the name Mathilde previously used in the subject position in a sentence with constituents nor in a sentence with no constituents, then it is strongly systematic w.r.t to how it can generalize to the novel use of the name Mathilde.
+On the other hand, if the system can only process this novel sentence if it has seen Mathilde being used in the same syntactic position, for example in the sentence Mathilde knows Olav, then it is not strongly systematic.
 Systematic generalization in embedded sentences.
-In this work, we will target a relaxed version of strong systematicity.
+In this work, we focus on a relaxed version of strong systematicity as defined in Definition 3.
 We relax requirement (ii) from the original definition so that we only work with samples that have embedded sentences.
 As we show in the following section, this makes it easier to quantify the level of systematicity required for a given generalization problem, which is the main focus of this work.
-Furthermore, we also relax (ii) so that it allows component parts in the test set to occur at the same syntactic position in some embedded sentence in the train set, but not in an embedded sentence that occur at the same position in the overall sequence.
+Furthermore, we also relax (ii) so that it allows component parts of the test set to occur at the same syntactic position in some embedded sentence in the train set, but not in an embedded sentence that occurs at the same position in the overall sequence.
 In order to compare models with and without compositional priors from the literature, we base our experiments on the SCAN grammar (Lake and Baroni, 2018).
@@
 3 Measuring Systematicity
-In this work, we will demonstrate that neural networks are able to generalize to embedded sentences where component parts occur at novel syntactic positions in the overall sequence.
-In our experiments, we will use a modified version of the SCAN grammar.
-We construct datasets where each dataset sample contains an input sequence $x\in X$ and an output sequence $y\in Y$.
-We discuss the specifics of this dataset generation procedure, including details about the context-free grammar (CFG) that generates $x$, in Section 3.2.
+In this work, we demonstrate that sequence-to-sequence models can generalize to embedded sentences where component parts occur at novel syntactic positions in the overall sequence.
+For our experiments, we construct datasets where each dataset sample contains an input sequence $x\in X$ and an output sequence $y\in Y$, based on a modified version of the SCAN grammar.
+We discuss the specifics of this dataset generation procedure, including details about the context-free grammar that generates $x$, in Section 3.5.
 In this section, we define a framework for measuring the degree of systematicity displayed by a systematic generalization problem.
-Throughout this paper, $x$ will comprise two embedded sentences conjoined by a conjunctive phrase.
-[^2] In contrast to the original SCAN, we do not have non-embedded sentences (e.g.
-sequences that do not have a conjunctive phrase).
+3.1 Task Description
+Throughout this paper, $x$ comprises two sentences conjoined by a conjunctive phrase.
+[^2] In contrast to the original SCAN, we do not have sentences without a conjunctive phrase.
 Consequently, $x$ can be modeled as a triple, with a conjunction $\operatorname{c}\in C$, and the two embedded sentences $e_{1},e_{2}\in\Sigma^{*}$ as arguments, $x=(e_{1},\operatorname{c},e_{2})$.
-The component parts of both embedded sentences are drawn from the same vocabulary $\Sigma=V\cup A$, where $V$ is a set of verbs and $A$ is a set of adverbial phrases, e.g: jump, run $\in V$, twice $\in A$, and and $\in C$.
+The component parts of both embedded sentences are drawn from the same vocabulary $\Sigma=V\cup A$, where $V$ is a set of verbs and $A$ is a set of adverbial phrases, e.g.: jump, run $\in V$, twice $\in A$, and and $\in C$.
 There exists a deterministic procedure $\varphi$ for transforming $x$ into $y$:
@@
 See Table 1 for more examples generated by our modified SCAN grammar.
-Since we are interested in systematic generalization, we need to also define the distribution of component parts in $e_{1}$ and $e_{2}$, and we are specifically interested in the case where the distribution of $e_{1}$ and $e_{2}$ is different in the training and test data, but they share a similar distribution when marginalized over position, which in this case refers to which side of the conjunctive phrase they appear in.
-In SCAN (Appendix A), $e_{1}$ and $e_{2}$ always contain one verb in $V$ and zero or more adverbial phrases from $A$ that control how many times that verb should be repeated, and in what direction.
+3.2 Component Distributions
+Since we are interested in systematic generalization, we also need to define the distribution of component parts in $e_{1}$ and $e_{2}$, and we are specifically interested in the case where the distribution of $e_{1}$ and $e_{2}$ is different in the training and test data, but they share a similar distribution when marginalized over position, which in this case refers to which side of the conjunctive phrase they appear in.
+In SCAN (Appendix A), $e_{1}$ and $e_{2}$ always contain one verb in $V$ and zero or more adverbial phrases from $A$ that control how many times that verb should be repeated, and in what direction it is executed.
 To create a systematic generalization problem that adheres to the definition given in Section 2.1, it is necessary to set the support of $e_{1}$ and $e_{2}$ to be different in the training and test distributions.
@@
 For $p^{\text{test}}$ we invert the distribution of $e_{1}$, making it a degenerate distribution with single-point support for the verb that is not in $e_{1}$ for $p^{\text{train}}$, while $e_{2}$ has the same distribution as in $p^{\text{train}}$.
-When generalizing from $p^{\text{train}}$ to $p^{\text{test}}$ in this scenario, a model will be exposed to all verbs in all permissible syntactic positions in some embedded sequence, but $v_{1}$ will never be seen in $e_{1}$ during training, while $e_{1}$ in $p^{\text{test}}$ always contains $v_{1}$.
+When generalizing from $p^{\text{train}}$ to $p^{\text{test}}$ in this scenario, a model will be exposed to all verbs in all permissible syntactic positions in some embedded sentence, but $v_{1}$ will never be seen in $e_{1}$ during training, while $e_{1}$ in $p^{\text{test}}$ always contains $v_{1}$.
 […]
 Figure 3: Example of distributions of verbs for $e_{1}$ (left) and $e_{2}$ (right).
-The main contribution of this work is to demonstrate that the difficulty of generalizing from $p^{\text{train}}$ to $p^{\text{test}}$ in this scenario depends on the entropy $\mathrm{H}$ of the distribution of verbs in $e_{2}$ for $p^{\text{train}}$:
+[…]
+Figure 4: Probabilities in the degenerate case for $e_{2}$
+3.3 Systematicity as Entropy
+The main contribution of this work is to demonstrate that the difficulty of generalizing from $p^{\text{train}}$ to $p^{\text{test}}$ in the scenario described above depends on the entropy $\mathrm{H}$ of the distribution of verbs in $e_{2}$ for $p^{\text{train}}$:
 $\mathrm{H}^{\text{train}}_{e_{2}}(\textrm{V})=-\sum_{v\in V}p^{\text{train}}_{e_{2}}(v)\log_{2}p^{\text{train}}_{e_{2}}(v),$
@@
 The degenerate case.
-[…]
-Figure 4: Probabilities in the degenerate case for $e_{2}$
 We note that $\mathrm{H}=0$ constitutes a special case in the described framework.
@@
 In this work, we experiment with two different approaches to increasing $H$, as illustrated in Figure 2.
-In the following sections we formalize these approaches, and in Section 4 we demonstrate how the performance of neural architectures scales for each type of increment.
+In the following sections, we formalize these approaches, and in Section 4 we demonstrate how the performance of neural architectures scales for each type of increment.
 3.4.1 Distribution Mixing
@@
 In our experiments, we use the standard conjunctive phrases in SCAN $|C|=2$.
-Furthermore, we let the two conjunctive phrases take the embedded sentences in the opposite order: if $v_{1}$ always occur in $e_{2}$ for $c_{1}$, then $v_{1}$ will always occur in $e_{1}$ for $c_{2}$.
+Furthermore, we let the two conjunctive phrases take the embedded sentences in the opposite order: if $v_{1}$ always occurs in $e_{2}$ for $c_{1}$, then $v_{1}$ will always occur in $e_{1}$ for $c_{2}$.
 4 Experiments
@@
 We also run a model with built-in compositional priors to showcase how a model that encodes verb equivariance can solve the task even when $\mathrm{H}$ is low.
-We have two main experimental settings, corresponding to the two methods for increasing $\mathrm{H}$ outlined in Section 3.1.1 and Section 3.1.2.
-We also conduct two supporting experiments.
+We have two main experimental settings, corresponding to the two methods for increasing $\mathrm{H}$ described in Section 3.4.1 and Section 3.4.2.
+We also conduct a supporting experiment to control for a potential confounding effect resulting from an increased sample size, and an experiment that estimates the influence of different position encoding types.
 For all experimental settings, $\mathrm{H}=0$ constitutes the most difficult generalization problem, and $\mathrm{H}=\log_{2}|V|=3$ constitutes the easiest problem.
@@
 The following sections provide details about our experimental setup.
-[^3] All code related to our experiments will be released upon publication at anonymous-url.
+[^3] Data generation scripts and code to reproduce our experiments can be found at https://github.com/ltgoslo/systematicity-entropy.
 4.1 Models
@@
 This group is used in the encoding of the input sequences in a way that pools the representation of any verb with the representation of the other verbs, even when they are not present in the input.
-This is a strong structural prior that makes the model able to solve the original SCAN splits efficiently.
+This is a strong structural prior that enables the model to solve the original SCAN splits efficiently.
 However, this approach is dataset-dependent, as the elements of the cyclic group must be defined manually.
@@
 For $p^{\text{train}}$, $e_{1}$ has a uniform distribution over $V\setminus\{v_{1}\}$, and $e_{2}$ is a mixture of a degenerate distribution at $v_{1}$ and a uniform distribution over $V\setminus\{v_{1}\}$, where we increase the total samples drawn from the uniform while decreasing the number drawn from the degenerate distribution.
-By doing this we increase $\mathrm{H}$ as described in section Section 3.1.1.
+By doing this, we increase $\mathrm{H}$ as described in section Section 3.4.1.
 We refer to this increase in $\mathrm{H}$ as vertical scaling, as the distribution over the verbs increases towards the uniform.
@@
 (b) Experiment 2: Horizontal scaling of entropy
-Figure 5: Results from the experiment 1 & 2. The accuracy and standard deviations are from five seeds.
+Figure 5: Results from the Experiments 1 & 2. The accuracy and standard deviations are from five seeds.
 4.3 Experiment 2: Horizontal Scaling
@@
 4.4 Effect of Number of Unique Samples
-As $H$ describes the relative distribution of verbs in $e_{2}$, it is possible to construct datasets of varying sizes that have the same $\mathrm{H}$.
+As $\mathrm{H}$ describes the relative distribution of verbs in $e_{2}$, it is possible to construct datasets of varying sizes that have the same $\mathrm{H}$.
 This can be achieved by increasing the total volume of unique samples while maintaining the same relative distribution.
@@
 In our experiments, there is an intuitive connection between the generalization from $p^{\text{train}}$ to $p^{\text{test}}$ and the ability to encode verb equivariance.
-As the verbs always occur on the first token position of $e_{1}$ and $e_{2}$, the positional information is central to both of our experiments.
-To assess the effect of positional information, we show how the choice of positional encoding scheme affects performance in the vertical scaling setting.
+This is exemplified by Gordon et al.
+(2019), which achieves high performance on the original SCAN tasks precisely by enforcing verb equivariance in the model architecture explicitly.
+As the verbs always occur on the first token position of $e_{1}$ and $e_{2}$, positional information is central to our experiments.
+To estimate the effect of positional information, we show how the choice of positional encoding scheme affects performance in the vertical scaling setting.
 In addition to the original absolute position embeddings from Vaswani (2017), we experiment with embeddings that encode relative positional information through disentangled attention (He et al., 2021) and rotary position embeddings (RoPE; Su et al., 2024).
@@
 4.6 Results and Discussion
+00111122223333000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111EntropyAccuracy300040006000
+(a) Effect of unique sample size on the Transformer
+00111122223333000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111EntropyAccuracyRotaryRelativeAbsol.
+(b) Effect of positional embedding type for the Transformer
+Figure 6: The result of the supplemental experiments. The accuracy and standard deviations are from five seeds.
 The results of Experiment 1 and Experiment 2 can be found in Figure 5.
@@
 This is not the case for the RNN and CNN, where we observe the opposite: The increase in support that increases $\mathrm{H}$ from 1 to 2 in Experiment 2 gives a higher performance increase than what the same increase in $\mathrm{H}$ does for Experiment 1.
-We argue that these results indicate that the Transformer is more information efficient than the RNN and CNN under full support, but that the models are similar under lacking support.
-In practice, this means that the Transformer is able to generalize to novel uses of component parts, as long as the overall syntactic pattern—constituted by the conjunctive phrase in our experiments—is frequent in the training data.
-This interpretation aligns with Loula et al.
+We argue that these results indicate that the Transformer is more information-efficient than the RNN and CNN under full support, but that the models are similar under lacking support.
+In practice, this means that the Transformer is quicker to generalize to novel uses of component parts, as long as the overall syntactic pattern—constituted by the conjunctive phrase in our experiments—is frequent in the training data.
+This interpretation aligns with the findings of Loula et al.
 (2018), showcasing how our framework can provide a theoretical explanation for previous results.
@@
 $\mathrm{H}=1$.
-This indicates a rather high lower bound on the entropy required for systematic generalization, both for vertical and horizontal scaling.
-This shows that systematic generalization remains an open problem when there is no strict architectural prior, and it remains unclear whether or not it is possible to build a model that performs well on low entropy without such priors.
+This indicates a rather high lower bound on the entropy required for systematic generalization, both for vertical and horizontal scaling, and it shows that all models are unable to meet requirement iii from Definition 3.
+Furthermore, our results show that systematic generalization remains an open problem when there is no prior, and it remains unclear whether or not it is possible to build a model that performs well on low entropy without architectural priors as in Gordon et al.
+(2019) or priors in the training procedure as in Lake and Baroni (2023).
 4.6.1 Supplemental results
-00111122223333000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111EntropyAccuracy300040006000
-(a) Effect of unique sample size on the Transformer
-00111122223333000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111EntropyAccuracyRotaryRelativeAbsol.
-(b) Effect of positional embedding type for the Transformer
-Figure 6: The result of the supplemental experiments. The accuracy and standard deviations are from five seeds.
 In Figure 6(a), we show the results of varying the number of unique samples at each level of $\mathrm{H}$.
@@
 We attribute the efficiency of absolute encodings to the fact that the positional information of the verbs is independent of the remainder of the sequence.
-This invariance has to be learned by the other embedding types, making the absolute variant more information efficient.
+This invariance has to be learned by the other embedding types, making the absolute variant more information-efficient.
 5 Previous Work
@@
 (2020).
-Most similar to our work is Zhou et al.
-(2023), who also studies how aspects of the SCAN grammar affects compositional generalization.
-We see compositionality as a set of primitives composed by a composition function, and our work focuses on how models capture this composition function depending on properties of the primitives.
-Their work, however, do not make this distinction, as their complexification process modify both the composition function and the distribution of primitives.
+The work most similar to ours is Zhou et al.
+(2023), who also studies how data-centric aspects affect compositional generalization.
+In our work, we see compositionality as characterized by a set of primitives composed by a composition function, and we focus on how models capture this composition function depending on properties of the primitives.
+Their work, however, does not make this distinction, as their complexification process modify both the composition function and the distribution of primitives.
 The main distinction is hence that our work relates compositionality to dataset-level regularities, while they relate it to sample-level features.
+Our work is also similar to Keysers et al.
+(2020), who quantifies the level of systematic generalization using distributional information.
+Specifically, they quantify the compositionality of an experiment as the divergence between the distribution of component parts and combinations in the training and test data.
+Their work differs from ours since they focus on the case where the distribution of component parts is as similar as possible between the training and test data, while our work focuses on the case where the distribution of component parts is inverted from the training to the test set for some parts of the input.
 In another domain, Wiedemer et al.
@@
 The authors show that the support of these variables must be the same in the training and test distributions for any model to generalize systematically.
-We take a more fine-grained approach, and our results demonstrate that support alone is insufficient as a predictor of systematic generalization; the performance also depends on the entropy, and higher levels of entropy in the training distribution can compensate for lower support.
-There is little existing research on how to measure systematicity from a data-centric perspective.
-However, there is ongoing work on formalizing compositionality in terms of data complexity (Elmoznino et al., 2024) and kernel-theory (Lippl and Stachenfeld, 2024).
-Most similar to our approach is Keysers et al.
-(2020), which also quantifies the level of systematic generalization using distributional information.
-Specifically, they quantify the compositionality of an experiment as the divergence between the distribution of component parts and combinations in the training and test data.
-Their work differs from ours since they focus on the case where the distribution of component parts is as similar as possible between the training and test data, while our work focuses on the case where the distribution of component parts is inverted from the training to the test set for some parts of the input.
+We take a more fine-grained approach and our results demonstrate that support alone is insufficient as a predictor of systematic generalization; the performance also depends on the entropy, and higher levels of entropy in the training distribution can compensate for lower support.
+There is also recent work on formalizing compositionality in terms of data complexity (Elmoznino et al., 2024), kernel-theory (Lippl and Stachenfeld, 2025), using graph formalisms (Ram et al., 2024), and decomposition strategies (Jarvis et al., 2023).
+Our work is similar to these in the sense that we try to formalize some properties of systematic generalization, which is closely tied to compositionality.
 6 Conclusion
-In this work, we show that the performance of sequence-to-sequence models scales with entropy.
-We also show that this scaling is independent of the number of unique samples.
-Contrary to previous work, our findings demonstrate that these models are capable of systematic generalization.
-However, this requires high entropy.
-This poses a new question: Is it possible to achieve systematicity also at low entropy without encoding explicit architectural priors?
-Through our formalizations, we hope to facilitate future work that tries to answer this question, providing a clear method for assessing progress in systematic generalization research.
+In this work, we show that systematic generalization of sequence-to-sequence models scales with information entropy.
+Through supporting experiments, we also show that this relationship is independent of the number of unique samples seen during training.
+Contrary to previous work, our findings demonstrate that these models are capable of systematic generalization even without any built-in architectural priors that incentivize compositional solutions.
+However, this requires that the training data have high entropy w.r.t.
+the parts that the target composition function operates on.
+This poses a new question: Is it possible to achieve systematicity also at low entropy without compositional priors?
+Through our formalizations, we hope to facilitate future work that attempts to answer this question, providing a clear method for assessing progress in systematic generalization research.
 Limitations
 Scope of our findings.
-In this work we have considered an aspect of systematicity that is concerned with embedded sentences as defined by Hadley (1994).
+In this work, we have considered an aspect of systematicity that is concerned with embedded sentences as defined by Hadley (1994).
 There is no guarantee that our findings generalize to scenarios that are different from ours, but that still falls in under some definition of systematic generalization.
@@
 Consequently, our method is constrained to cases where fine-grained distributional information is obtainable.
+Acknowledgements
+We acknowledge Sigma2—the National Infrastructure for High-Performance Computing and Data Storage in Norway—for providing access to the LUMI super-computer, part of the EuroHPC Joint Undertaking, hosted by CSC (Finland) and the LUMI consortium.
+We also want to thank Lilja Øvrelid, Erik Velldal, and David Samuel at the Language Technology Group for discussions and feedback on the manuscript.
 Appendix A Modified SCAN grammar
@@
   • Transformer sweep: 120 hours.
-  • Permutation equivariant sweep: 24 hours
+  • Permutation equivariant sweep: 24 hours.
   • RNN training run: 600 hours.